# Super-Higher-Form Symmetries

### v1

@@ 1 Introduction
The possibility to define a supersymmetric theory on a supermanifold [^1] For an introduction to supersymmetric theories in supermanifolds, see Fre:2017bwd ; Castellani:2023tip ., or more simply on a superspace, which is the local version of a supermanifold, allows for accomplishing this program in a very efficient way. In fact, several ingredients discussed in the aforementioned literature can be easily translated into a supersymmetric framework. First of all, differential calculus in supermanifolds V1 ; V2 ; V3 ; V4 ; Belopolsky:1997bg ; Belopolsky:1997jz ; Voro , leads to the notion of superdifferential forms $J^{(p|q)}$ that locally can be expressed in terms of fundamental $(1|0)$-forms $dx$’s and $d\theta$’s related to the even and odd coordinates of the superspace, with the coefficients of the expansion being ordinary superfields. In a $(n|m)$-dimensional supermanifold, they are featured by a form number $0\leq p\leq n$ and a so-called picture number $0\leq q\leq m$. Superdifferential forms satisfying the closure condition $dJ^{(p|q)}=0$ are the natural candidates for defining super-higher-form currents, that is, the generalization of higher-form conserved currents to a supersymmetric set-up. The general theory of integration on supermanifolds Witten:2012bg allows then to define the corresponding generators as integrals of the currents on $(p|q)$-dimensional super-hypersurfaces $\Sigma_{(p|q)}$ embedded into the supermanifold, $Q(\Sigma_{(p|q)})=\int_{\Sigma_{(p|q)}}\!\!J^{(p|q)}$. Thanks to the closure of the super differential form, these are topological operators that generate a $(n-p-1|m-q)$-higher-form symmetry. Charged objects can be constructed in general, which are operators localized on a $(n-p-1|m-q)$—hypersurface.
@@ 2.1 Maxwell theory in $n$ dimensions
$S^{\prime}+S_{min}^{\prime}$ $=$ $S+S_{min}+\int_{\mathcal{M}}\left(d\lambda^{(1)}\wedge D^{(4)}-\frac{1}{2}d\lambda^{(1)}\wedge d\lambda^{(1)}\wedge B^{(2)}_{CW}\right)\,.$ (2.16)
@@ 3 Super-Higher-Form Symmetries
We consider a field theory defined in the supermanifold $\mathcal{SM}$, with dimension $(n|m)$, where $n$ is the bosonic dimension and $m$ is the fermionic one. Locally, it is homeomorphic to a superspace $\mathbb{R}^{(n|m)}$ parametrized by coordinates $z^{A}=(x^{a},\theta^{\alpha})$, where $a=0,\dots,n-1,\alpha=1,\dots,m$. Flat supervielbeins are given by $V^{a}=dx^{a}+\theta^{\alpha}\gamma^{a}_{\alpha\beta}d\theta^{\beta}$, $\psi^{\alpha}=d\theta^{\alpha}$, for a suitable choice of Clifford gamma matrices. The superdifferential $d$ is defined as $d=V^{a}\partial_{a}+\psi^{\alpha}D_{\alpha}$, where $D_{\alpha}$ are the supercovariant derivatives satisfying $\{D_{\alpha},D_{\beta}\}=-\gamma^{a}_{\alpha\beta}\partial_{a}$. From the identity $dV^{a}=\psi\gamma^{a}\psi$ it follows $d^{2}=0$.
@@
Ordinary vector superform currents in superspace correspond to $J^{(n-1|m)}\equiv\star J^{(1|0)}$, whose expansion in given in (3.2) for $p=0$. For $p=1$, the closure of $J^{(n-2|m)}$ gives rise to the standard conservation of the U(1) supercurrent in superspace, $D_{\alpha}(J^{\alpha}+\gamma_{a}^{\alpha\beta}D_{\beta}J^{a})=0$ Cremonini:2020zyn . This is the lowest representative of a chain of generalized conserved supercurrents obtained by increasing the form degree $p$ as in the ordinary bosonic case and/or changing the picture number.
@@
It follows that on a supermanifold without boundaries, if $\star J^{(p+1|m)},\star J^{(p+1|0)}$ are closed forms (higher-superfom conserved currents), the integrals in (3.4) define topological quantities. They correspond to the generators of a $(p|m)$-integral form and a $(p|0)$-superform symmetry, respectively, whose parameters $\lambda^{(p|q)}$ are now an integral form if $q=m$ or a superform if $q=0$.
@@ 4 Super-Maxwell theory
Looking for operators charged under the action of gCW symmetries generated by (4.2), we now prove that the supermonopole $\mathcal{D}(\Sigma_{(1|4)})$ introduced above is also charged respect to them. To this end, we first couple the gCW currents (4.61) to background gauge fields, which amounts to consider in the path integral the generalized action
@@ 6 Conclusions
Lastly, we have not addressed the case of pseudo-differential forms $J^{(p|q)}$ when $q<m$, but these forms could potentially lead to new interesting conserved currents. They may be associated with super-hypersurfaces that have a reduced number of fermionic coordinates. However, the mathematics involved in the construction of these objects is considerably more complex, as the vector spaces required to describe them are naturally infinite-dimensional. Although there have been some new efforts to incorporate them into the discussion, further mathematical tools still need to be developed.
@@ Appendix A Superforms and Integral Forms
Given a supermanifold ${\mathcal{M}}^{(n|m)}$, parametrized by the coordinates $z^{A}=(x^{a},\theta^{\alpha})$ where $a=0,\dots,n-1$ and $\alpha=1,\dots,m$, we can construct the exterior algebra either from the canonical 1-form basis $dx^{a},d\theta^{\alpha}$ or from the anolonomy basis $V^{a},\psi^{\alpha}$ (a.k.a. supervielbeins). The differential is given by
@@
where $D_{\alpha}$ is the covariant derivative.
@@
In addition, we can introduce the contraction operator
@@
where we adopted the conventional notation used for Cartan calculus where $D_{\alpha}$ is the vector field along which we compute the contraction $\iota_{\alpha}$ acting on superforms (A.2). If we denote by $\iota_{a}=\iota_{\partial_{a}}$ the conventional contraction along the coordinate vector field $\partial_{a}$ we have
@@
The difference between $\iota_{\alpha}$ and $\iota_{a}$ is
$\iota_{a}\iota_{b}=-\iota_{b}\iota_{a}\,,~{}~{}~{}~{}~{}\iota_{\alpha}\iota_{\beta}=\iota_{\beta}\iota_{\alpha}\,.~{}~{}~{}$ (A.9)
@@
The integral forms form a differential complex which is isomorphically dual to the complex of superforms and the notion of the Hodge dual can be established
@@
and the details are given in Castellani:2015ata ; Castellani:2015dis ; Castellani:2023tip ; Castellani:2016ezd . The relevant feature of the integral forms is the they can be integrated on suitable super-cycles (hyper surfaces embedded into a supermanifold). In this way, we can obtain a well-defined expression

### v2

@@ 1 Introduction
The possibility to define a supersymmetric theory on a supermanifold [^1] For an introduction to supersymmetric theories in supermanifolds, see Fre:2017bwd ; Castellani:2023tip ., or more simply on a superspace, which is the local version of a supermanifold, allows for accomplishing this program in a very efficient way. In fact, several ingredients discussed in the aforementioned literature can be easily translated into a supersymmetric framework. First of all, differential calculus in supermanifolds BL1 ; BL2 ; Manin ; V1 ; V2 ; V3 ; V4 ; Belopolsky:1997bg ; Belopolsky:1997jz ; Voro , leads to the notion of superdifferential forms $J^{(p|q)}$ that locally can be expressed in terms of fundamental $(1|0)$-forms $dx$’s and $d\theta$’s related to the even and odd coordinates of the superspace, with the coefficients of the expansion being ordinary superfields. In a $(n|m)$-dimensional supermanifold, they are featured by a form number $0\leq p\leq n$ and a so-called picture number $0\leq q\leq m$. Superdifferential forms satisfying the closure condition $dJ^{(p|q)}=0$ are the natural candidates for defining super-higher-form currents, that is, the generalization of higher-form conserved currents to a supersymmetric set-up. The general theory of integration on supermanifolds Witten:2012bg allows then to define the corresponding generators as integrals of the currents on $(p|q)$-dimensional super-hypersurfaces $\Sigma_{(p|q)}$ embedded into the supermanifold, $Q(\Sigma_{(p|q)})=\int_{\Sigma_{(p|q)}}\!\!J^{(p|q)}$. Thanks to the closure of the super differential form, these are topological operators that generate a $(n-p-1|m-q)$-higher-form symmetry. Charged objects can be constructed in general, which are operators localized on a $(n-p-1|m-q)$—hypersurface.
@@ 2.1 Maxwell theory in $n$ dimensions
$S^{\prime}+S_{min}^{\prime}$ $=$ $S+S_{min}+\int_{\mathcal{M}}\left(d\lambda^{(1)}\wedge B_{CW}^{(4)}-\frac{1}{2}d\lambda^{(1)}\wedge d\lambda^{(1)}\wedge B^{(2)}_{CW}\right)\,.$ (2.16)
@@ 3 Super-Higher-Form Symmetries
We consider a field theory defined in the supermanifold $\mathcal{SM}$, with dimension $(n|m)$, where $n$ is the bosonic dimension and $m$ is the fermionic one. Locally, it is isomorphic to a superspace $\mathbb{R}^{(n|m)}$ parametrized by coordinates $z^{A}=(x^{a},\theta^{\alpha})$, where $a=0,\dots,n-1,\alpha=1,\dots,m$. Flat supervielbeins are given by $V^{a}=dx^{a}+\theta^{\alpha}\gamma^{a}_{\alpha\beta}d\theta^{\beta}$, $\psi^{\alpha}=d\theta^{\alpha}$, for a suitable choice of Clifford gamma matrices. The superdifferential $d$ is defined as $d=V^{a}\partial_{a}+\psi^{\alpha}D_{\alpha}$, where $D_{\alpha}$ are the supercovariant derivatives satisfying $\{D_{\alpha},D_{\beta}\}=-\gamma^{a}_{\alpha\beta}\partial_{a}$. From the identity $dV^{a}=\psi\gamma^{a}\psi$ it follows $d^{2}=0$.
@@
Ordinary vector superform currents in superspace correspond to $J^{(n-1|m)}\equiv\star J^{(1|0)}$, whose expansion is given in (3.2) for $p=0$. For $p=1$, the closure of $J^{(n-2|m)}$ gives rise to the standard conservation of the U(1) supercurrent in superspace, $D_{\alpha}(J^{\alpha}+\gamma_{a}^{\alpha\beta}D_{\beta}J^{a})=0$ Cremonini:2020zyn . This is the lowest representative of a chain of generalized conserved supercurrents obtained by increasing the form degree $p$ as in the ordinary bosonic case and/or changing the picture number.
@@
It follows that on a supermanifold without boundary, if $\star J^{(p+1|m)},\star J^{(p+1|0)}$ are closed forms (higher-superfom conserved currents), the integrals in (3.4) define topological quantities. They correspond to the generators of a $(p|m)$-integral form and a $(p|0)$-superform symmetry, respectively, whose parameters $\lambda^{(p|q)}$ are now an integral form if $q=m$ or a superform if $q=0$.
@@ 4 Super-Maxwell theory
Looking for operators charged under the action of gCW symmetries generated by (4.2), we now prove that the supermonopole $\mathcal{D}(\Sigma_{(1|4)})$ introduced above is also charged respect to them. To this end, we first couple the gCW currents (4.61) to background gauge fields, which amounts to considering in the path integral the generalized action
@@ 6 Conclusions
Lastly, we have not addressed the case of Voronov-Zorich superforms $J^{(p|q)}$ with $q<m$, though these forms could potentially lead to new interesting conserved currents. They may be associated with super-hypersurfaces that have a reduced number of fermionic coordinates. However, the mathematics involved in the construction of these objects is considerably more complex, as the vector spaces required to describe them are naturally infinite-dimensional. Although there have been some new efforts to incorporate them into the discussion, further mathematical tools still need to be developed.
@@ Appendix A Superforms and Integral Forms
Given a supermanifold ${\mathcal{M}}^{(n|m)}$, parametrized by the coordinates $z^{A}=(x^{a},\theta^{\alpha})$ where $a=0,\dots,n-1$ and $\alpha=1,\dots,m$, we can construct the exterior algebra either from the canonical 1-form basis $dx^{a},d\theta^{\alpha}$ or from the anholonomic basis $V^{a},\psi^{\alpha}$ (a.k.a. supervielbeins). The differential is given by
@@
where $D_{\alpha}$ is the covariant derivative, while $dx^{a}$ and $d\theta^{\alpha}$ are anticommuting and commuting differentials, respectively.
@@
As in conventional geometry, we can introduce the contraction operator $\iota_{X}$ along a vector field $X$ on the manifold, acting on the exterior algebra. The contraction operator $\iota_{X}$ reduces the form degree by one unit. In the present context, we can define the contraction operator along the even-coordinate vector field $\frac{\partial}{\partial x^{a}}$ and the one along the odd-coordinate vector field $\frac{\partial}{\partial\theta^{\alpha}}$. However, it is more convenient to introduce covariant contractions. Along the covariant vector field $D_{\alpha}$ we define
@@
whereas along bosonic vectors $\partial_{a}\equiv\frac{\partial}{\partial x^{a}}$ we define
$\iota_{a}\equiv\iota_{\partial_{a}}=\frac{\partial}{\partial V^{a}}\,,$ (A.8)
They act on superforms (A.2) and are dual to the supervielbein $(V^{a},\psi^{\alpha})$,
@@
The commutation relation among $\iota_{\alpha}$ and $\iota_{a}$ are
$\iota_{a}\iota_{b}=-\iota_{b}\iota_{a}\,,~{}~{}~{}~{}~{}\iota_{\alpha}\iota_{\beta}=\iota_{\beta}\iota_{\alpha}\,,~{}~{}~{}\iota_{\alpha}\iota_{b}=\iota_{b}\iota_{\alpha}\,.~{}~{}~{}$ (A.10)
@@
The integral forms form a differential complex which is isomorphically dual to the complex of superforms. The notion of the Hodge dual can be established
@@
and the details are given in Castellani:2015ata ; Castellani:2015dis ; Castellani:2023tip ; Castellani:2016ezd . In our work, we have considered only two special cases, namely those with $q=0$ and $q=m$. For these two cases we verified the isomorphism, which easily follows from the fact that they are finite dimensional spaces. For the generic case, the isomorphism might require a deeper analysis. In that respect, we also mention the related work VZ .
The relevant feature of the integral forms is the they can be integrated on suitable super-cycles (hyper surfaces embedded into a supermanifold). In this way, we can obtain a well-defined expression